I am trying to change this into slope-intercept form. The equation is 2x + 5y = 12. I divided by 2 first which turns the equation x + 5y = 6. I then divided 5 on both sides and turned the equation to x + y = 6/5. I then multiplied x on both sides and got y = 6/5x. Our teacher tiled us that we don't have to graph if b/y-intercept is a fraction. Did I solve this equation correctly?
Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

I am trying to change this into slope-intercept form.
The equation is 2x + 5y = 12.
I divided by 2 first which turns the equation x + 5y = 6.
You didn't divide the 5 by 2.
-----
I then divided 5 on both sides and turned the equation to x + y = 6/5.
You didn't divide the x by 5.
-------
I then multiplied x on both sides and got y = 6/5x.
It's not clear what you did there.
-------------
Our teacher tiled us that we don't have to graph if b/y-intercept is a fraction. Did I solve this equation correctly?
==============================
2x + 5y = 12
Isolate y, get it in the form y = mx + b
Subtract 2x
5y = -2x + 12
Divide by 5
y = (-2/5)x + 12/5

Answer by richwmiller(17219)

No
your way with many mistakes
2x + 5y = 12
divide by 2
x + 5y/2 = 6
divide by 5
x/5+y/2=6/5
the right way
2x + 5y = 12
get just y on one side
subtract 2x from both sides
5y=12-2x
divide by 5
y=12/5-2x/5
rearrange
y=-2/5x+12/5


another way
2x + 5y = 12
slope is -a/b
slope is -2/5
y intercept is y when x=0
2*0 + 5y = 12
5y=12
y=12/5
now we have the slope and the y intercept
y=-2/5x+12/5